questions:
1. what is the term recession refers to?
2. what is the most volatile component of aggregate demand?
3. Assume that the Marginal propensity to consume(MPC) is 0.75, and the investment spending rises by $25 billion. By how much will the real GDP change?
4. What is the equation for the velocity f money?
5. if Namibians suddenly decide to hold more cash for carrying on transactions and for precautionary reasons, what is more likely to happen?
6. Supply-side fiscal policy is generally enacted through what?
7. The Phillips curve describes the relationship between what and what?
8. Phillips curve is an illustration of what?
9. In the crude equation of exchange where MV=PY, what is likely to happen?
10. A bill of exchange is a money market financial security that is traded where?
11. What is the most important conclusions from the keynesian liquidity preference theory?
12. Where changes in the growth of money supply in an economy can originate from?
1
Expert's answer
2014-09-18T13:06:00-0400
1. what is the term recession refers to?
In economics, a recession is a business cycle contraction. It is a general slowdown in economic activity. Macroeconomic indicators such as GDP (gross domestic product), investment spending, capacity utilization, household income, business profits, and inflation fall, while bankruptcies and the unemployment rate rise.
2. what is the most volatile component of aggregate demand?
The most volatile component of aggregate demand is investment.
3. Assume that the Marginal propensity to consume(MPC) is 0.75, and the investment spending rises by $25 billion. By how much will the real GDP change?
S = s*Y, s = 1 - c, so change in Y = 25/(1 - 0.75) = $100 billion.
4. What is the equation for the velocity f money?
MV = PQ, so velocity V = PQ/M, where P - price, Q - quantity, M - money base.
5. if Namibians suddenly decide to hold more cash for carrying on transactions and for precautionary reasons, increase in the inflation rate is more likely to happen.
6. Supply-side fiscal policy is generally enacted through an increase in government spending.
7. The Phillips curve describes the relationship between inflation and unemployment.
8. Phillips curve is an illustration of an inverse correlation between an economy's unemployment rate and inflation rate
9. In the crude equation of exchange where MV=PY, mistakes in real money base calculation are likely to happen.
10. A bill of exchange is a money market financial security that is usually traded internationally.
11. What is the most important conclusions from the keynesian liquidity preference theory?
The LIQUIDITY PREFERENCE THEORY is the idea that investors demand a premium for securities with longer maturities, which entail greater risk, because they would prefer to hold cash, which entails less risk. The more liquid an investment, the easier it is to sell quickly for its full value. Because interest rates are more volatile in the short term, the premium on short- versus medium-term securities will be greater than the premium on medium- versus long-term securities.
12. Changes in the growth of money supply in an economy can originate from change in the reserve requirement rate.